# Hybrid deep learning and iterative methods for accelerated solutions of viscous incompressible flow

### v1

@@ 2.2 Decoupled immersed boundary projection method
$\nabla\cdot\mathbf{u}$ $=$ $0$ (11)
@@
[…]
@@ 2.3 HyDEA: Hybrid Deep lEarning line-search directions and iterative methods for Accelerated solutions
DLSM employs a deep neural network to provide the mapping between the $r_{k}$ and $e_{k}$, with the algorithmic implementation formally described in Algorithm 1.111Note that the DCDM of Kaneda et al. [32] represents a special case of DLSM, where the line-search directions are made additionally $M$-orthogonal by a GS process. We shall demonstrate in Section 3.4 that the GS orthogonalizaiton is unnecessary, as a properly trained neural network already suggests effective line-search directions.
@@
The workflow of HyDEA is presented in Fig. 1, where it receives an initial solution and its corresponding residual for the current CFD time step. Subsequently, it selects to perform whether the DLSM or a CG-type method for the current iteration. After each iteration, the $L2$-norm of the updated residual is evaluated and if it falls below a predefined threshold, the iterative process terminates. Otherwise, the procedure repeats, progressively refining the solution until the convergence criterion is met.
Figure 1: The workflow of HyDEA, $n$ denotes the time step for an unsteady CFD simulation, $k$ represents the iteration index for solving the discrete PPE at the current time step, and $atol$ is the predefined absolute tolerance. The initial value for iteration is taken from last time step: $\delta p_{k}=\delta p^{n-1}$
Further details of HyDEA are illustrated in Fig. 2. As which solver executed first is potentially essential, HyDEA offers two distinct choices as indicated by the red-dashed circle in Fig. 2. The two implementations in Fig. 2 and Fig. 2 correspond to the two versions of HyDEA for one round of the hybrid algorithm: the former represents a CG-type method as the initial solver, while the latter employs DLSM as the initial solver. Moreover, $k$ represents the current iteration number, and $mod$ is the modulo operation. For the two respective methods ${DLSM}_{count}$ and ${CG-type}_{count}$ record the numbers of consecutive iterations while $Num_{DLSM}$ and $Num_{CG-type}$ denote the predefined maximum consecutive iterations.
@@ 2.4.1 Dataset construction
The critical step in dataset construction is to generate a set of approximate eigenvectors $Q_{m}=[\mathbf{q}_{0},\mathbf{q}_{1},\cdots,\mathbf{q}_{m-1}]\in R^{a\times m}$ (where $m<a$), which can effectively represent the entire spectrum of $M$. Specifically, the Lanczos iteration first transforms $M$ to a low-dimensional tridiagonal matrix $T_{m}\in R^{m\times m}$ [50], yielding the relation
@@
where $V_{m}\in R^{a\times m}$ is the orthogonal matrix whose columns consist of the orthogonal vectors generated during the Lanczos iteration. Subsequently, the eigenvalues of $T_{m}$ form a diagonal matrix $\Lambda_{m}\in R^{m\times m}$ with non-decreasing eigenvalues on the diagonal referred to as approximate eigenvalues or Ritz values of $M$. Moreover, the eigenvectors of $T_{m}$ form a matrix $U_{m}\in R^{m\times m}$. Thereafter, $Q_{m}=V_{m}U_{m}$ forms the approximate eigenvectors of $M$, commonly known as Ritz vectors. Finally, the dataset is constructed as follows
$r^{i}$ $=$ $\frac{\sum_{j=0}^{m-1}c_{j}^{i}\mathbf{q}_{j}}{||\sum_{j=0}^{m-1}c_{j}^{i}\mathbf{q}_{j}||_{2}},$ (40)
@@ 2.4.2 Deep operator network and model training
The mapping between $e_{k}$ and $r_{k}$ in Eq. (38) can be understood as an operator, which we employ the DeepONet to learn its biased version governed by the dataset. The unstacked version of DeepONet employed here consists of a branch network and a trunk network [16] as illustrated in Fig. 3, where the branch network is a U-Net architecture made of convolutional neural networks (CNN) [51] concatenated by a feedforward neural network (FNN) and the trunk network is a simple FNN.
@@
$[bs,c,h,w]$ is the feature map, where $bs$ is $batchsize$, $c$ represents the number of channels, $h$ and $w$ are the sizes of the input in two spatial directions. In U-Net, $T$, $2\cdot T$, $4\cdot T$, etc., represent the numbers of feature map channels at specific network layers and $T=40$ by default. $A$, $B$ and $C$ are the number of neurons in FNNs and are assigned as $100$, $200$ and $100$ by default, respectively. A detailed analysis of these parameters are presented in B.
@@ 2.5 Technical details
super-resolution $640\times 640\downarrow$ $512\times 512\downarrow$
@@ 3.1.1 $Re=1000$
The geometric configuration and boundary conditions of the flow are illustrated in Fig. 4. The computational domain is a square with side length $H_{x}=H_{y}=H=1$. The x-component of the velocity for the top lid is set to be $u=1$ and all other velocity components of the boundaries are 00. The kinematic viscosity $\nu$ is $0.001$, and the CFD time step  $\Delta t=0.004$. Therefore, Reynolds number $Re=uH/\nu=1000$. The computational domain is discretized using a uniform Cartesian grid with resolution of $128\times 128$. Consequently, the size of the input feature map for the branch network of DeepONet is set to $h=128$ and $w=128$.
@@
To prepare for the training dataset, we adopt the parameter values $m=3000$ and $b=0.6$. A systematic sensitivity analysis of their influence on HyDEA’s performance is presented in C. Furthermore, we set $Num_{CG-type}=3$ and $Num_{DLSM}=2$ as the maximum executive iterations of the two respective solvers in each round of the hybrid algorithm and remove a detailed analysis on the two values into D. The termination residual expressed in $L2$-norm for the iterative calculation is set to $\epsilon=10^{-6}$, being consistent with the official benchmark of $\mathtt{PetIBM}$.
@@
Figure 5: The iterative residuals of solving the PPE at the $10th$, $100th$ and $1000th$ time steps for a 2D lid-driven cavity flow at $Re=1000$. (a)-(c) HyDEA (CG+DLSM-1). (d)-(f) HyDEA (ICPCG+DLSM-1). (g)-(i) HyDEA (JPCG+DLSM-1). (j)-(l) HyDEA (MGPCG+DLSM-1).
@@
Acceleration ratio ×$\times 3.83$$\times 1.88$$\times 3.56$$\times 1.64$
@@
Some comments are in order; the reduction of computational time in Table 2 is less impressive than the reduced number of iterations in Fig. 5. This performance discrepancy stems from multiple factors. Foremost, HyDEA requires frequently bidirectional data tensor transfers between GPU (neural networks) and CPU (iterative methods), leading to significant execution delays due to communications. Secondly, the DLSM method is implemented in $\mathtt{Python}$, which inherently limits its computational efficiency due to the interpreted nature of the language. Additionally, the current network architecture may be optimized for faster inference. While these technical factors limit the current computational efficiency of HyDEA, they also represent a clear opportunity for future optimization. We emphasize that the present work primarily focuses on establishing the algorithmic framework of HyDEA and examine its generalization capability.
@@ 3.1.2 $Re=3200$
To prepare for the training dataset, we adopt the parameter values $m=7000$ and $b=0.6$. A systematic sensitivity analysis of their influence on HyDEA’s performance is also provided in C. Furthermore, we set $Num_{CG-type}=3$ and $Num_{DLSM}=2$ as the maximum executive iterations of the two respective solvers in each round of the hybrid algorithm. The termination criterion is again with $\epsilon=10^{-6}$.
@@
Figure 8: The iterative residuals of solving the PPE at the $10th$, $100th$ and $1000th$ time steps for 2D lid-driven cavity flow at $Re=3200$. (a)-(c) HyDEA (CG+DLSM-2). (d)-(f) HyDEA (ICPCG+DLSM-2).
@@
Table 3: The wall-time acceleration ratio and computational time of solving the PPE by HyDEA for a duration of $10,000\Delta t$ at $Re=3200$.
@@
Acceleration ratio ×$\times 5.96$$\times 2.82$$\times 5.64$$\times 2.08$
@@ 3.2 Case 2: One stationary circular cylinder immersed in 2D lid-driven cavity flow
Figure 17: The iterative residuals of solving the PPE for 2D lid-driven cavity flow with an immersed stationary circular cylinder at $Re=3200$ by ICPCG, HyDEA (ICPCG+DLSM-2) and HyDEA (CG+DLSM-2) . (a) $10th$ time step. (b) $100th$ time step. (c) $1000th$ time step.
@@ 3.3 Case 3: Two stationary elliptical cylinders immersed in 2D lid-driven cavity flow
Figure 21: The iterative residuals of solving the PPE for 2D lid-driven cavity flow with two immersed stationary elliptical cylinders at $Re=3200$ by ICPCG, HyDEA (ICPCG+DLSM-2), and HyDEA (CG+DLSM-2). (a) $10th$ time step. (b) $100th$ time step. (c) $1000th$ time step.
@@ 3.4 Case 4: 2D flow around an inline oscillating cylinder at $Re=100$
The DCDM of Kaneda et al. [32] represents a special case of DLSM, where the line-search directions are made additionally $M$-orthogonal by a GS orthogonalization process. We evaluate respective performance of DLSM module and DCDM module within the HyDEA framework. We compare iterative residuals of HyDEA (ICPCG+DLSM-2) and HyDEA (ICPCG+DCDM-2) for solving the PPE at the $10th$, $1000th$ and $3400th$ time steps in Fig. 29. The results show that HyDEA achieves similar convergence rates when equipped with either DLSM-2 or DCDM-2, suggesting that GS orthogonalization of DeepONet-predicted line-search directions provides negligible benefits.

### v2

@@ 2.2 Decoupled immersed boundary projection method
$\nabla\cdot\mathbf{u}$ $=$ $0.$ (11)
@@
$\mathbf{f}(\mathbf{s},t)$ $=$ $\int_{\Gamma}\mathbf{F}(\mathbf{s},t)\delta(\mathbf{x}-\mathbf{X}(\mathbf{s},t))d\mathbf{s},$ (12)
$\mathbf{U}(\mathbf{s},t)$ $=$ $\int_{\Omega}\mathbf{u}(\mathbf{x},t)\delta(\mathbf{X}(\mathbf{s},t)-\mathbf{x})d\mathbf{x},$ (13)
@@ 2.3 HyDEA: Hybrid Deep lEarning line-search directions and iterative methods for Accelerated solutions
DLSM employs a deep neural network to provide the mapping between the $r_{k}$ and $e_{k}$, with the algorithmic implementation formally described in Algorithm 1. The DCDM of Kaneda et al. [32] represents another example of a line-search method, where the line-search directions are made additionally $M$-orthogonal by a GS process. In the present work, both DLSM and DCDM employ identical neural network architecture, training dataset and training methodology. We shall demonstrate in Section 3.4 that the GS orthogonalization is unnecessary, as a properly trained neural network already suggests effective line-search directions.
@@
The workflow of HyDEA is presented in Fig. 1, where it receives an initial solution and its corresponding residual for the current CFD time step. Subsequently, it selects to perform whether the DLSM or a CG-type method for the current iteration. After each iteration, the $L2$-norm of the updated residual is evaluated and if it falls below a predefined threshold, the iterative process terminates. Otherwise, the procedure repeats, progressively refining the solution until the convergence criterion is met. Notably, when the procedure switches from DLSM to a CG-type method, the CG-type method’s iterative state is completely restarted from scratch, with the DLSM’s final solution serving as the initial guess for the subsequent iterations.
Figure 1: The workflow of HyDEA, $n$ denotes the time step for an unsteady CFD simulation, $k$ represents the iteration index for solving the discrete PPE at the current time step, and atolatol is the predefined absolute tolerance. The initial value for iteration is taken from last time step: $\delta p_{k}=\delta p^{n-1}$.
Further details of HyDEA are illustrated in Fig. 2. As which solver executed first is potentially essential, HyDEA offers two distinct choices as indicated by the red-dashed circle in Fig. 2. The two implementations in Fig. 2 and Fig. 2 correspond to the two versions of HyDEA for one round of the hybrid algorithm: the former represents a CG-type method as the initial solver, while the latter employs DLSM as the initial solver. Moreover, $k$ represents the current iteration number, and modmod is the modulo operation. For the two respective methods ${DLSM}_{count}$ and ${CG-type}_{count}$ record the numbers of consecutive iterations while $Num_{DLSM}$ and $Num_{CG-type}$ denote the predefined maximum consecutive iterations.
@@ 2.4.1 Dataset construction
The critical step in dataset construction is to generate a set of approximate eigenvectors $Q_{m}=[\mathbf{q}_{0},\mathbf{q}_{1},\cdots,\mathbf{q}_{m-1}]\in\mathbf{R}^{a\times m}$ (where $m<a$), which can effectively represent the entire spectrum of $M$. Specifically, the Lanczos iteration first transforms $M$ to a low-dimensional tridiagonal matrix $T_{m}\in\mathbf{R}^{m\times m}$ [50], yielding the relation
@@
where $V_{m}\in\mathbf{R}^{a\times m}$ is the orthogonal matrix whose columns consist of the orthogonal vectors generated during the Lanczos iteration. Subsequently, the eigenvalues of $T_{m}$ form a diagonal matrix $\Lambda_{m}\in\mathbf{R}^{m\times m}$ with non-decreasing eigenvalues on the diagonal referred to as approximate eigenvalues or Ritz values of $M$. Moreover, the eigenvectors of $T_{m}$ form a matrix $U_{m}\in\mathbf{R}^{m\times m}$. Thereafter, $Q_{m}=V_{m}U_{m}$ forms the approximate eigenvectors of $M$, commonly known as Ritz vectors. Finally, the dataset is constructed as follows
[…]
@@ 2.4.2 Deep operator network and model training
The mapping between $e_{k}$ and $r_{k}$ in Eq. (38) can be understood as an operator, which we employ the DeepONet to learn its biased version governed by the dataset. The unstacked version of DeepONet employed here consists of a branch network and a trunk network [16] as illustrated in Fig. 3. The computational domain is discretized using a uniform Cartesian grid. This representation allows the flow field to be naturally treated as an image, where each grid point corresponds to a pixel, with the values of $r_{k}$ at different grid points forming the pixel values. The U-Net architecture made of convolutional neural networks (CNN), known for its exceptional performance in computer vision [51], is particularly suitable for processing such image-structured data. Therefore, the branch network employs a U-Net concatenated by a feedforward neural network (FNN). The trunk network applies a simple FNN.
@@
$[bs,c,h,w]$ is the feature map, where $bs$ is batchsizebatchsize, $c$ represents the number of channels, $h$ and $w$ are the sizes of the input in two spatial directions. In U-Net, $T$, $2\cdot T$, $4\cdot T$, etc., represent the numbers of feature map channels at specific network layers and $T=40$ by default. $A$, $B$ and $C$ are the number of neurons in FNNs and are assigned as $100$, $200$ and $100$ by default, respectively. $K$ represents the depth of the U-Net architecture. A detailed analysis of these parameters are presented in B.
@@ 2.5 Technical details
Super-resolution $640\times 640\downarrow$ - $512\times 512\downarrow$ -
@@
The code and data are available on GitHub at https://github.com/HMB9666/HyDEA.
2.6 Error Analysis of HyDEA
Prior to applying HyDEA to CFD, we perform an error analysis by solving a one-dimensional(1D) Poisson equation with a known, frequency-controllable solution:
$\frac{d^{2}u}{dx^{2}}$ $=$ $f(x),\quad x\in\Omega=[-1,1]$ (43)
$u(x)$ $=$ $g,\quad x\in\partial\Omega$ (44)
where $u(x)$ denotes the field function to be computed. The source term $f(x)$ and the boundary values $g$ are prescribed based on the manufactured solution $u_{exact}$:
$u_{exact}=\sin(x)-\frac{1}{4}\sin(4x)+\frac{1}{8}\sin(8x)-\frac{1}{16}\sin(16x)+\frac{1}{24}\sin(24x)$-
$\frac{1}{48}\sin(48x)+\frac{1}{128}\sin(128x)-\frac{1}{256}\sin(256x).$ (45)
A finite difference discretization is applied to Eq. (43) on a uniform grid of 2000 nodes, utilizing a second-order central difference scheme for the Laplace operator. The resulting linear equation is solved using both the CG method and HyDEA (CG + DLSM) from a zero initial guess, with the termination residual expressed in $L2$-norm set to $\epsilon=10^{-7}$. To enable a clear observation of the distinct error evolution behaviors of both the CG and DLSM components within HyDEA, $Num_{CG-type}$ and $Num_{DLSM}$ are both set to 10. The neural network architecture employs the DeepONet as illustrated in Fig. 3. The U-Net depth $K$ is set to 7, with the input feature map, coordinates, and convolutional kernels all adapted for 1D configuration. To prepare for the training dataset, we adopt the parameter values $m=800$ and $b=0.6$. All other hyperparameters and training settings remain consistent with those detailed in Section 2.4.2.
The iterative residuals of the 1D Poisson equation using both the CG method and HyDEA(CG+DLSM) are presented in Fig. 4. HyDEA receives the same initial residual (shown in solid circle) as in CG, then executes $Num_{CG-type}=10$ CG iterations (solid pentagon), followed by $Num_{DLSM}=10$ DLSM iterations (empty pentagon), and continues this alternating pattern. In total, HyDEA reaches the predefined tolerance by 7 rounds of hybrid algorithm, outperforming the CG method alone. Furthermore, it can be observed that the iterative residual reduction of HyDEA in the initial-stage iteration is primarily driven by the DLSM component, while the CG component becomes the dominant contributor in the later-stage iteration. The accuracy of HyDEA is confirmed in Fig. 5 by the close match between the numerical solution and $u_{exact}$.
Figure 4: Iterative residuals of solving the 1D Poisson equation by CG and HyDEA (CG+DLSM).
Figure 5: Comparison of the numerical solution obtained by HyDEA (CG+DLSM) and $u_{exact}$.
Furthermore, the evolution of the iterative error’s power spectral density ($PSD(e)$) for HyDEA (CG+DLSM) as a color map is shown in Fig. 6, while the corresponding evolution profiles at wavenumbers $\kappa=1$, $10$ and $50$ are presented in Fig. 6. The results reveal that the DLSM component in HyDEA acts as the primary driver for eliminating low-wavenumber error energy during the initial-stage iteration. Nevertheless, in the later-stage iteration, DLSM excites growth and oscillations in high-wavenumber error energy, which is subsequently eliminated effectively after switching to the CG method. This observation is consistent with the evolution of iterative residuals shown in Fig. 4, which implies that the complementary behavior of CG and DLSM in HyDEA leads to a rapid decay of the iterative residuals.
Figure 6: Evolution of the $PSD(e)$ for HyDEA (CG+DLSM). (a) Color map. (b) Evolution profiles at $\kappa=1$, $10$ and $50$, with gray areas marking the intervals of DLSM activity.
@@ 3.1.1 $Re=1000$
The geometric configuration and boundary conditions of the flow are illustrated in Fig. 7. The computational domain is a square with side length $H_{x}=H_{y}=H=1$. The x-component of the velocity for the top lid is set to be $u=1$ and all other velocity components of the boundaries are 0. The kinematic viscosity $\nu$ is $0.001$, and the CFD time step  $\Delta t=0.004$. Therefore, Reynolds number $Re=uH/\nu=1000$. The computational domain is discretized using a uniform Cartesian grid with resolution of $128\times 128$. Consequently, the size of the input feature map for the branch network of DeepONet is set to $h=128$ and $w=128$. The U-Net depth $K$ is set to 5 for this case and all subsequent sections, as it provides satisfactory empirical performance, with no further parameter analysis presented.
@@
The training dataset is prepared with the parameter values $m=3000$ and $b=0.6$, which required 268 seconds for construction and 18 hours for model training. A systematic sensitivity analysis of their influence on HyDEA’s performance is presented in C. Furthermore, we set $Num_{CG-type}=3$ and $Num_{DLSM}=2$ as the maximum executive iterations of the two respective solvers in each round of the hybrid algorithm and remove a detailed analysis on the two values into D. The termination residual expressed in $L2$-norm for the iterative calculation is set to $\epsilon=10^{-6}$, being consistent with the official benchmark of $\mathtt{PetIBM}$. The influence of this parameter on the performance of HyDEA is discussed in E.
@@
Figure 8: Iterative residuals of solving the PPE at the $10th$, $100th$ and $1000th$ time steps for 2D lid-driven cavity flow at $Re=1000$. (a)-(c) HyDEA (CG+DLSM-1). (d)-(f) HyDEA (ICPCG+DLSM-1). (g)-(i) HyDEA (JPCG+DLSM-1). (j)-(l) HyDEA (MGPCG+DLSM-1).
@@
Acceleration ratio $\times 3.83$ $\times 1.88$ $\times 3.56$ $\times 1.64$
@@
Computational time of
DLSM-1 (s)
59 57 60 54
Computational time of
DeepONet (s)
39 38 39 35
GPU-CPU transfer
time (s)
4.62 4.48 4.67 3.93
Some comments are in order; the reduction of computational time in Table 2 is less impressive than the reduced number of iterations in Fig. 8. This performance discrepancy stems from multiple factors. Foremost, the DLSM component in HyDEA is implemented in $\mathtt{Python}$ and $\mathtt{PyTorch}$. The interpreted nature of the language, combined with the complexity of the neural network architecture, thereby limiting the overall computational efficiency, as clearly demonstrated in Table 2. Additionally, the bidirectional tensor transfers between the GPU (neural networks) and CPU (iterative methods) introduce minor communication overhead. While these technical factors limit the current computational efficiency of HyDEA, they also represent a clear opportunity for future optimization. We emphasize that the present work primarily focuses on establishing the algorithmic framework of HyDEA and examine its generalization capability.
@@ 3.1.2 $Re=3200$
The training dataset is prepared with the parameter values $m=7000$ and $b=0.6$, which required 3138 seconds for construction and 41 hours for model training. A systematic sensitivity analysis of their influence on HyDEA’s performance is also provided in C. Furthermore, we set $Num_{CG-type}=3$ and $Num_{DLSM}=2$ as the maximum executive iterations of the two respective solvers in each round of the hybrid algorithm. The termination criterion is again with $\epsilon=10^{-6}$.
@@
Figure 11: Iterative residuals of solving the PPE at the $10th$, $100th$ and $1000th$ time steps for 2D lid-driven cavity flow at $Re=3200$. (a)-(c) HyDEA (CG+DLSM-2). (d)-(f) HyDEA (ICPCG+DLSM-2).
@@
Table 3: The wall-time acceleration ratio and computational time for solving the PPE by HyDEA for a duration of $10,000\Delta t$ at $Re=3200$.
@@
Acceleration ratio $\times 5.96$ $\times 2.82$ $\times 5.64$ $\times 2.08$
@@
Computational time of
DLSM-2 (s)
100 103 102 97
Computational time of
DeepONet (s)
62 63 63 60
GPU-CPU transfer
time (s)
9.56 8.69 9.71 8.54
@@ 3.2 Case 2: One stationary circular cylinder immersed in 2D lid-driven cavity flow
Figure 20: Iterative residuals of solving the PPE for 2D lid-driven cavity flow with an immersed stationary circular cylinder at $Re=3200$ by ICPCG, HyDEA (ICPCG+DLSM-2) and HyDEA (CG+DLSM-2) . (a) $10th$ time step. (b) $100th$ time step. (c) $1000th$ time step.
@@ 3.3 Case 3: Two stationary elliptical cylinders immersed in 2D lid-driven cavity flow
Figure 24: Iterative residuals of solving the PPE for 2D lid-driven cavity flow with two immersed stationary elliptical cylinders at $Re=3200$ by ICPCG, HyDEA (ICPCG+DLSM-2), and HyDEA (CG+DLSM-2). (a) $10th$ time step. (b) $100th$ time step. (c) $1000th$ time step.
@@ 3.4 Case 4: 2D flow around an inline oscillating cylinder at $Re=100$
The DCDM of Kaneda et al. [32] represents a special case of DLSM, where the line-search directions are made additionally $M$-orthogonal by a GS orthogonalization process. We evaluate respective performance of DLSM module and DCDM module within the HyDEA framework. We compare the iterative residuals of HyDEA (ICPCG+DLSM-2) and HyDEA (ICPCG+DCDM-2) for solving the PPE at the $10th$, $1000th$ and $3400th$ time steps in Fig. 32. The results show that HyDEA achieves similar convergence rates when equipped with either DLSM-2 or DCDM-2, suggesting that GS orthogonalization of DeepONet-predicted line-search directions provides negligible benefits.